OpenStudy (anonymous):

the length and the breadth of a square are increased by 40% and 30% resp.the area of resulting rectangle exceeds the area of square by

OpenStudy (anonymous):

@kropot72 ,@ganeshie8 ,@Compassionate

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@mathslover

OpenStudy (mathstudent55):

You have a square, so the length and width are equal. Start with an easy number for the side, such as 10. What is the area of a square with side 10?

OpenStudy (anonymous):

A1 = L x L A2 = 1.4L x 1.3 L calculate either A2-A1 or, better, A2/A1

OpenStudy (anonymous):

l^2-1.82l^2=.82l^2

OpenStudy (mathstudent55):

The area of a square with side 10 is 10 * 10 = 100 Now increase the length by 40% to get 14. Increase the width by 30% to get 13. Now you have a rectangle with length and width 14 and 13 units, respectively. the area of this rectangle is 14 * 13= 182

OpenStudy (anonymous):

100-182/100*100

OpenStudy (anonymous):

82%

OpenStudy (mathstudent55):

Let's find the percent increase from 100 to 182. \(percent~change = \dfrac{182 - 100}{100} \times 100 = 82\%\)

OpenStudy (anonymous):

yes, 1.4 x 1.3 = 1.82 an 82% increase

OpenStudy (mathstudent55):

Correct. The percent increase is 82%. The increase in area is 1.82 times.

OpenStudy (anonymous):

@mathstudent55 is showing the best way to do one of these problems if the numbers aren't given, meaning whatever number you pick will work a great trick to learn